From GlueXWiki
Jump to: navigation, search

November 29, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Package #1 subsystems
    • Package #2 production
    • Package #4 testing (Vlad)
  2. Corrosion tests (Vlad)
  3. Engineering
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 (Lubomir)
    • Third package wire resolution FDC E-log
    • Plans for testing packages #1 and #4
  6. Other


Participants: Eugene, Dave, Chris, Nick, Vlad, Simon, Beni, Glenn, and Lubomir.


- Dave: now the techs are back and we are working on package #2. Tina and Mike are cutting/gluing new foils; Anatoly tested already the six new foil sets that we have received. Anatoly started gluing the daughter cards on the cathodes that were done before. Priority is to finish the two type-3 cathodes that will be needed in any package configuration.

- On package #1 Vlad installed the ground connections before Thanksgiving, then this week we installed the cooling loops and yesterday Chris and Nick started working on the HV cabling. The package will be ready at the beginning of the next week. We decided not to install the cooling manifolds so far: we don't need them now and in case we need to open the package we will not be able to re-use the cooling tubes. On package #3 it turned out the splittings of the cooling tubes are in wrong places due to misinterpreting the drawings. Correcting this will take ~2 days but will do it later.

- Package #4 was assembled (the two missing cells) just before Thanksgiving and flushed with gas over the holidays: now oxygen is 151ppm. Vlad tested all the cells: currents on cells from #6 to #2: 21/-3, 44/-4, 42/-3, 50/-3, 110/-100 nA. Three cathode channels with bad contacts between the daughter cards and the pre-amps found; after removing/inserting the pre-amps they start working.

- On cell#1 package #4:HV could be applied only to several HV sectors +/-HV1, +/-HV4, and +HV3. On the rest we found shorts between the HV and the lower cathode and therefore we opened it today: as expected we found a broken wire, field 47. It is not clear when and how the wire broke. Certainly it was in place when the wire frame was installed in the package, but later after closing the package somehow it broke in the middle, the only explanation so far that it may have had a kink in this place (Casey). Also some wire frame deformation must be involved, say even on the pins it stayed deformed but after tightening the package got flat and the wire broke? The tension of the wire was 122g very close to its neighbors. Casey is replacing already the wire; the plan is to finish it by tomorrow so that we can put the package together before the weekend.

Cathode corrosion test

- Vlad: test #3 (Ar/CO2 gas, at ~70degC). Two samples failed before Thanksgiving (14 days from the start of the tests): one EPDM and the other dummy, i.e. no O-ring on the sample, just a strip with two wires soldered to the ends. The second EPDM sample is OK already 23 days. We assume it was the solder that failed again; it failed before at 100degC in four days. From these results one estimates about 0.5eV activation energy, which means this type of soldering on 2 micron copper would fail after ~190 days at room temperature, good we don't solder on 2 micron copper! As for the Viton+Apiezon the highest estimate so far is that it will survive for at least ~1300 days.


- Without Bill not much to say.


- Chris: the ground clips were shipped will be here soon. Finally the daughter cards arrived and we started using them at Blue Crab. However, it turned out they have not been tested. Chris/Nick will take them next week for testing, except the cards that were put on one cathode that will be tested at Blue Crab.

Tests at EEL126

- The third package was moved from 126 to Blue Crab before Thanksgiving. Before that Lubomir collected ~5 days data with the special three-cell-cathode configuration (as explained before in Elog #96) needed to extract the time-to-distance function and estimate the wire resolution. The new results with higher statistics and better track quality are shown in entry #97: the polynomial parametrization of this function is given there; the average wire resolution is estimated to be ~230 microns. This is an upper estimate since the cathode resolution is expected to be worse than the one estimated from the pickle-fence histogram.

- The A/C units for 126 were delivered today and the the plan is to start the installation on Monday. Tomorrow the techs will be there to clean the place: shelves, tool boxes etc.

- Related to that we decided to postpone the move of the first package for testing to 126. First, there will be installation work there, second, we expect the HVAC to be down for the Christmas holidays as it was last year. We will still have time to test package #1 in January and package #4 in February. By that time we will be ready with package #2.

- Beni plans to re-adjust the range of all mass flow controllers. The dynamic range is a factor of ten. A balance has to made if we want to be able to use gas mixtures in the range from 40/60 to 90/10 Ar/CO2 and a spare factor in case we want to increaser the flow to compensate for eventual gas contamination. After the meeting Beni and Lubomir decided: 1) to be able to change the gas mixture from 40/60 to 90/10, 2) have 250 sccpm nominal flow per package, 3) to have the possibility to increase the flow by a factor of 1.6 from the nominal, 4) to have identical flow controllers for Ar and CO2. The safety factor of 1.6 may seem low, but the point is that we know the gas contamination is now low, it is defined by the sealing technology used, and we don't expect significant deterioration.